Find the z score that has 80% of its area to the right? Show work
Answer #1

solution

Using standard normal table,

P(Z > z) =80 %

= 1 - P(Z < z) = 0.80

= P(Z < z ) = 1 - 0.80

= P(Z < z ) = 0.20

= P(Z <- 0.84 ) = 0.20  

z- score = - 0.84 (using standard normal (Z) table )